OpenStudy (anonymous):

A. –3 B. 0 C. 3 D. 8

OpenStudy (anonymous):

whats the question

OpenStudy (anonymous):

its the link

OpenStudy (alexandervonhumboldt2):

3/8 * 8/b=1 8/b=8/3 1/b=8/3*1/8 1/b=1/3 b=3

OpenStudy (anonymous):

dang... you go glen coco